OpenStudy (anonymous):

write an equation of a line in slope-intercept form for that is parallel to y=5x-10 and passes through the point (9,-3)

OpenStudy (shinalcantara):

since the line is parallel to the equation y=5x-10, then the line's slope is the same as that of the line of the equation. REMEMBER that y-intercept form is the defined by the formula y=mx + b where m is the slope and b being the y-intercept. Based on the given equation, what would be the slope of the line?

OpenStudy (anonymous):

the answer is y=5x-48 can you please explain how that is the answer @shinalcantara

OpenStudy (shinalcantara):

since the slope of the line is 5 and you have a point (9,-3), all you have to remember when dealing with equation of the line with given slope and a point is the slope formula \[m = \frac{ y2-y1 }{ x2-x1 }\] substitute the values \[ 5 = \frac{ y - (-3) }{ x-9 }\] cross multiply (x-9) \[5(x-9)=y+3\] \[5x - 45 = y + 3\] rearranging \[y = 5x - 45 - 3\] \[y = 5x - 48\]